OpenStudy (anonymous):

i need help 1.(-mn8)2 2.(3x3y)2(-4x2y4)3

OpenStudy (unklerhaukus):

\[(-mn^8)^2=(-mn^8)(-mn^8)=m^2n^{16}\]

OpenStudy (anonymous):

2012Sorching summer comes, in order to thank everyone, characteristic, novel style, varieties, low price and good quality, and the low sale price. Thank everyone ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== $33 True Religion jeans, Ed Hardy jeans,LV,Coogi jeans,Affliction jeans $30 Air Jordan shoes,Shox shoes,Gucci,LV shoes 50%Discount winter fashion :Sandle,t-shirt,caps,jerseys,handbag and brand watches!!! $15 Ed Hardy ,LV ,Gucci Bikini $15 Polo, Ed Hardy, Gucci, LV, Lacoste T-shirts $25 Coach,Gucci,LV,Prada,Juicy,Chanel handbag, $10 Gucci,Ed Hardy sunglasses $9 New Era caps. give you the unexpected harvest ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.scnshop.com ) =====

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

\[(3x^3y)^2(-4x^2y^4)^3=(3x^3y)(3x^3y)(-4x^2y^4)(-4x^2y^4)(-4x^2y^4)\]

OpenStudy (anonymous):

how whould i do that one with the y?

OpenStudy (unklerhaukus):

\[(3x^3y)(3x^3y)(−4x^2y^4)(−4x^2y^4)(−4x^2y^4)\] \[=(9x^6y^2)(-12x^6y^{12})\]

OpenStudy (unklerhaukus):

can you do the last step?